OpenStudy (anonymous):

Find any rational roots of P(x). P(x)=3x^4+2x^3-9x^2+4

OpenStudy (anonymous):

try 1, -1 first then 2, -2 then 4,-4 if these don't work we have to move on to fractions

OpenStudy (anonymous):

there is more than 1 so even if they work there wtill might be a fraction. Oh and with this i have to show my work and do it the textbook way not by trial and error

OpenStudy (anonymous):

So what exactly is the textbook way for your textbook

OpenStudy (anonymous):

there is no other way to do it you have to check. if that is "trial and error' that is what you need to do

OpenStudy (anonymous):

Well, in practicality my teacher told me to graph the function.. and then you can use synthetic division for the ones that aren't "exact"

OpenStudy (anonymous):

But unless you have some sort of sixth sense of math you really don't have a set method.. unless your textbook does, in which case i'd want to know.

OpenStudy (anonymous):

ya thats the way i am just having trouble relating it to this specifc problem

OpenStudy (anonymous):

1 works by your eyeballs because if \[P(x)=3x^4+2x^3-9x^2+4\] then \[P(1)=3+2-9+4=0\]

OpenStudy (anonymous):

well for this one i would guess -2

OpenStudy (anonymous):

but its just a guess

OpenStudy (anonymous):

add up the coefficients and see that you get 0 then you can factor as \[(x-1)\times (something)\] and then you can either factor the "something" or else use the quadratic formula

OpenStudy (anonymous):

you only need to find one zero, then factor and end up with a quadratic, and you can always find the zeros of any quadratic

OpenStudy (anonymous):

scratch that, this is a polynomial of degree 4, so you factor and end up with a cubic when you do that, you will see that 1 is a zero of the cubic polynomial as well, so you will factor again as \((x-1)(x-1)(something)\)

OpenStudy (anonymous):

x+2

OpenStudy (anonymous):

but there really isn't an exact method to it is all.